TreatMeNot.Inc is a U.S.-based pharmaceutical company that specializes on flu vaccines. In the past two months they have been working on developing a vaccine against coronavirus. Drug development was performed in cooperation with Rambam R&D Lab (the Lab). After a series of tests, completed yesterday, the vaccine was proven effective. The Lab has charged the company $12M for their services. The first payment of $6M was made 3 months ago, and the remaining $6M is due tomorrow.

The company’s analysts anticipate generating $100,000M (that is, $100B) in vaccine sales in the first year which will grow at an annual negative growth rate of sixty percent (that is, g=-60%) for the subsequent 7 years. The costs of production are estimated to be 72% of sales. To be conservative in their estimates, the analysts of TreatMeNot.Inc have decided to ignore cash flows starting from year 9 and onward, assuming that at that point the disease will be fully eradicated (in other words, the project lasts for eight years in total).

If the company decides to produce the vaccine, it will have to purchase vaccine manufacturing equipment right away, which will cost $150M. The equipment will be depreciated using straight-line method to the residual value of $25M over its ten-year life span. It can be sold at the end of the project (that is, at the end of year 8), for $20M to Pfizer, Inc. The production will take place in a building that is currently used for nursing students training program (the company offers the program as part of its portfolio of products and services). The program, which generates pre-tax profits of $5M a year, will be suspended for the duration of the project’s life but will resume after the end of the project (and can potentially last forever afterwards).

The company’s corporate tax rate is 20 percent; the required return is 12 percent, and the company is all-equity financed. In answering the question, assume that all the cash flows and the discount rate are nominal. Unless specified otherwise, all the cash flows are pre-tax, and arrive/are paid at the end of the specified period. If necessary, you can also assume that if EBIT in a given year is negative, the company receives a tax refund of 20% from the government (the same year).

  1. Should the company produce the vaccine? To receive full credit for the question, show the time line, the FCF, and all of your calculations. Answers in Excel are acceptable.
  2. How would your answer change if you assume that revenue growth could range between -40% and -80%? Use sensitivity/scenario analysis to address the question. I encourage you to use Excel for this part to save time. After performing the calculations, explain in one-two sentences how this additional information would change your overall conclusions.
  3. Johnson & Johnson has approached TreatMeNot.Inc this morning with an offer to buy the vaccine patent. If the deal goes through, TreatMeNot.Inc will not be able to manufacture the vaccine. What is the fair price at which TreatMeNot.Inc should accept the offer? Explain your answer in 1-2 sentences.
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Required Return =12%=0.12
N=Year   of Cash Flow
$12 million charged by Lab is sunk cost and not relevant for this decision
A Cost of Equipment($ million) $150
B Residual Value($ million) $25
C=(A-B)/10 Annual depreciation $12.50
Book Value at end of 8 years=150-(8*12.5)= $50.00 million
Salvage Value at end of 8 years $20 million
Loss on Salvage $30 million
Tax Savings on Loss=30*20% $6 million
After tax salvage cash flow at end of 8 years $26 million (20+6)
Sales Revenue in Year1($ million) $100,000
Sales Revenue in Year2($ million) (100000*(1-0.6)
Sales Revenue in Year(N+1)=0.4*(Sales Revenue in Year (N)
YEARWISE CASH FLOWS($ million)
N Year 0 1 2 3 4 5 6 7 8
A Initial Cash Flow in Equipment -$150
b Annual Revenue $100,000 $40,000 $16,000 $6,400 $2,560 $1,024 $410 $164
c=b*72% Annual Cost of production(Before Depreciation) -$72,000.0 -$28,800.0 -$11,520.0 -$4,608.0 -$1,843.2 -$737.3 -$294.9 -$118.0
d Depreciation Expenses -$12.5 -$12.5 -$12.5 -$12.5 -$12.5 -$12.5 -$12.5 -$12.5
e Reduction of current pre tax profit -$5.0 -$5.0 -$5.0 -$5.0 -$5.0 -$5.0 -$5.0 -$5.0
f=b+c+d+e Before tax Operating Income $27,982.5 $11,182.5 $4,462.5 $1,774.5 $699.3 $269.2 $97.2 $28.4
g=f*(1-0.2) After Tax Operating Income $22,386.0 $8,946.0 $3,570.0 $1,419.6 $559.4 $215.4 $77.8 $22.7
h Add : Depreciation (non cash expense) $12.5 $12.5 $12.5 $12.5 $12.5 $12.5 $12.5 $12.5
P=f+g Annual Operating Cash Flow $22,398.5 $8,958.5 $3,582.5 $1,432.1 $571.9 $227.9 $90.3 $35.2
S Terminal Salvage Case $26
CF=A+P+S Net Cash Flow -$150 $22,398.5 $8,958.5 $3,582.5 $1,432.1 $571.9 $227.9 $90.3 $61.2 SUM
PV=CF/(1.12^N) Present Valure -$150 $19,998.7 $7,141.7 $2,550.0 $910.1 $324.5 $115.4 $40.8 $24.7 $30,955.9
NPV=Sum of PV Net Present Value(NPV) $30,955.9
Company should produce the vaccine
NPV is Positive

NET PRESENT VALUE IF REVENUE GROWTH IS -40% 4 5 -$150 c=b72% YEARWISE CASH FLOWS($ million) Year Initial Cash Flow in EquipmeNET PRESENT VALUE IF REVENUE GROWTH IS -80% 4 5 6 А -$150 $6.4 -$4.6 $0.0 -$5.0 YEARWISE CASH FLOWS($ million) Year Initial C